Acetylene gas is used in industrial works such as welding metals. The


temperature needed to weld is approximately 2200°C at 1500 psia. Calculate its


density at this given temperature and pressure in kgm per cubic meter

1
Expert's answer
2022-02-17T04:36:31-0500

"pV=\frac mMRT,"

"\rho=\frac{pM}{RT},"

"\rho=13~\frac{kg}{m^3}."